RI Ra R4

Consider the circuit shown above. Assume the battery is ideal and there is no current through the switch when it is open.

With the switch open, if all four resistances are equal, rank order the resistors in terms of the power dissipated in them. In other words match "1" with the resistor with the largest power, "2" with the next largest, etc.

If any two or more resistors have the same power, match them with the same number (both "1", or both "2", etc.)
Answer #1

if switch is open,

Req = R + 2R*R/3R = (5/3)*R

if battar have V volt,

current in circult I = V/Req = V/(5/3)*R = (3/5)*(V/R) = (3/5)*I0

power dissipated = I2R

here all ressistance is equal (R ohm) than power depands on current.

so current in R1 = (3/5)*I0

current in R2 = (3/5)*I0*(2R)/(3R) = (2/5)*I0

current in R3 and R4 = (1/5)*I0

so Power in R1 > Power in R2 > Power in R3 = Power in R4